OpenStudy (aldana.navea):

Need Help with this one: Solve for x y=(x-3)m

OpenStudy (michele_laino):

first step: please divide both sides by \(m\)

OpenStudy (michele_laino):

namely: \[\frac{y}{m} = \frac{{\left( {3 - x} \right)m}}{m}\] please simplify

OpenStudy (aldana.navea):

okay one sec (:

OpenStudy (aldana.navea):

so this is my answer so far, because we cross out the two m's right? \[\frac{ y }{ m }= (3-x)\]

OpenStudy (michele_laino):

that's right! now I multiply both sides by \(-1\), so I get: \[ - \frac{y}{m} = - 3 + x\]

OpenStudy (michele_laino):

finalli, I add \(3\) to both sides: \[ - \frac{y}{m} + 3 = - 3 + x + 3\] please simplify

OpenStudy (michele_laino):

finally*

OpenStudy (aldana.navea):

thank you (: give me one second

OpenStudy (michele_laino):

hint: we can write this: \[ - 3 + x + 3 = x + 3 - 3 = ...?\]

OpenStudy (aldana.navea):

would it be -9+x=x?

OpenStudy (michele_laino):

it is: \[x + 3 - 3 = x\] so we get: \[x = - \frac{y}{m} + 3\]

OpenStudy (michele_laino):

since: \(3-3=3+(-3)=0\)

OpenStudy (aldana.navea):

Oh okay! I added the 3's except of crossing them out, thank you!

OpenStudy (michele_laino):

:)